### v2

# Quantum mechanical closure of partial differential equations with symmetries

Abstract
We develop a framework for the dynamical closure of spatiotemporal dynamics governed by partial differential equations. Employing the mathematical framework of quantum mechanics to embed the original classical dynamics into an infinite dimensional quantum mechanical system, we use the space of quantum states to model the unresolved degrees of freedom of the original dynamics and the framework of quantum measurement to predict their contributions to the resolved dynamics. The embedded dynamics is projected to finite dimension by a positivity preserving discretization process. Based on methods from operator valued kernels and delay embedding, the compressed finite dimensional representation of the dynamics is invariant under the spatial symmetries of the original dynamics. We develop a data driven formulation of the scheme that can be realized numerically and apply it to a dynamical closure problem for the shallow water equations. The results demonstrate that our closure model can accurately predict the main features of the true dynamics, including for out of sample initial conditions.
keywords:
dynamical closure, parametrization, quantum mechanics, kernel methods, delay embedding, transfer operator, shallow water equations
[…]
[1]organization=Department of Mathematics, Dartmouth College, city=Hanover, state=NH 03755, country=USA
1 Introduction
Complex dynamical systems often encompass degrees of freedom that evolve over a wide range of temporal and spatial scales with their evolution depending on different physical regimes. A challenging problem in such multiscale, multiphysics systems is trying to include dynamical information for degrees of freedom that either are too computationally expensive to simulate directly or depend on fully or partially unknown physical processes [1].
The approach employed by closure or parametrization methods is to derive surrogate dynamical models that can be used to approximate the contributions of fine grain degrees of freedom to the dynamics of coarse grain processes we can simulate consistently. Early applications include turbulent flows and climate dynamics [2, 3, 4, 5], but closure methods have since found consequential applications in a wide range of fields [6]. For example, in atmospheric dynamics a technique known as super-parametrization employs deterministic or stochastic column models with explicitly resolved cloud microphysics as surrogate models of small scale moist convection, embedded within the discretization cells of a coarse atmospheric model [7, 8]. In plasma dynamics there have been efforts to derive models for subgrid scale processes in fluid models as well as to derive fluid closures of kinetic models, incorporating both physical modeling and data driven approaches [9, 10, 11, 12, 13, 14, 15, 16].
Related to closure is model reduction, which aims to derive models of reduced complexity that agree in a prescribed sense with the original dynamics. For instance, reduced models can be derived using operator theoretic methods that attempt to justify the model’s physical consistency in a statistical sense [17, 18, 19, 20, 21] or directly through computational methods [22, 23, 24, 25, 26, 27, 28, 29].
Mathematically, closure methods depend on a separation of the state space of the original dynamical system into two components: (1) the resolved degrees of freedom, which represent the coarse grain processes we can simulate or measure consistently and with tractable cost; (2) the unresolved degrees of freedom, which are considered too complex or expensive to simulate or measure directly. Once the unresolved degrees of freedom have been identified, one has to define a surrogate model for their fluxes—namely, their contributions—to the resolved dynamics. How to define such a model and whether such a model should be deterministic or stochastic are active areas of research that use various mathematical tools to offer alternative approaches [30, 31, 1, 32, 33].
One approach toward this problem is defining a surrogate model as a functional relationship between the resolved and unresolved degrees of freedom. In this way, the unresolved variables are determined by a deterministic function of the resolved variables. A widespread approach of defining such functions that was especially popular early on relies on expert knowledge of the modeled physical process to derive formulas that seek to predict the averaged fluxes to the resolved variables [34]. These formulas typically feature tunable parameters that are calibrated by analyzing numerical, observational or experimental data through regression or inference techniques [35, 36, 37, 38]. More recently, an alternative data driven approach has emerged where functional relationships are not derived based on expert physical reasoning but are discovered from data using machine learning methods [39, 40, 41, 42, 43, 27, 44, 45, 28, 46, 47, 48, 49].
Modeling the unresolved degrees of freedom by a deterministic function of the resolved degrees corresponds to a reduction in the dimension of the full state space of the original dynamics. Regardless of the complexity of the employed functional relationship, any dynamical independence the unresolved variables had has been relinquished. One important consequence of this reduction in state space dimension is that the resulting parametrized system may experience a loss in its ability to produce complex spatiotemporal behavior such as chaotic dynamics [50]. If that happens, the developed parametrized system will inevitably struggle to consistently approximate the original dynamics.
To prevent the reduction of the full state space to the resolved state space one can introduce a surrogate state space for the unresolved degrees of freedom and use that space to model their contributions to the resolved dynamics. One such approach has been to use stochastic methods to model the unresolved degrees of freedom, thereby complementing the resolved state space with the event space of the introduced stochastic variables [51, 52, 31, 53, 32, 54].
Independently of whether the unresolved variables are represented using a deterministic or stochastic framework, the main goal of closure methods is to construct a surrogate model that can accurately and consistently approximate the original dynamics. Additionally, it should do so with a tractable computational cost and without imposing a strong reduction in the complexity of the predicted dynamical behavior. Despite the wide variety of mathematical tools employed, reduction and closure methods can be studied within a unified operator theoretic framework. Often referred to as the Mori-Zwanzig formalism, the framework uses projection operators to describe the reduction of fine grain degrees of freedom to coarse grain ones, resulting in exact governing equations for the coarse grain variables [17, 18, 19, 21]. The derived governing equations offer useful physical interpretations of the terms that arise in the projection process, and can be used to derive error bounds and consistency proofs for downstream reduced models.
A framework called quantum mechanical closure (QMCl) [55] has been recently developed which models the unresolved degrees of freedom as a quantum mechanical system. This approach can be viewed as a generalization of surrogate modeling based on stochastic models. Building on earlier work on quantum mechanical data assimilation [56, 57], QMCl encodes statistical information about the unresolved degrees of freedom in a quantum density operator [58, 59, 60] that evolves under the transfer operator of the dynamical system [61]. The fluxes of the unresolved variables are represented by quantum observables, while the generation of the required flux terms in the course of timestepping is carried out using tools from the mathematical theory of quantum measurement [58, 59, 60]. The resulting closure scheme is expressed in terms of linear operators acting on Hilbert spaces. This enables automatic positivity preservation (important for modeling sign definite physical quantities such as density or temperature) as well as data driven approximation using kernel integral operator methods [62, 63, 64] with rigorous consistency guarantees in the limit of infinite training data. QMCl has thus far been applied to the closure of prototype chaotic ordinary differential equations such as the Lorenz 63 and multiscale Lorenz 96 models, where it has shown promising results for the reconstruction of time correlation functions and long term statistics of observables.
1.1 Main results
In this work we extend the QMCl framework to the case of spatiotemporal dynamics governed by partial differential equations. In the extended framework, information about the unresolved degrees of freedom is encoded in fields of quantum density operators which are embedded within the spatial discretization cells of the coarse model. This produces a framework that can naturally take advantage of the spatial structure exhibited by partial differential equations, such as dynamical symmetries. More specifically, we employ elements of the vector valued spectral analysis (VSA) framework introduced in [65] to develop a data driven scheme that can effectively and efficiently model the closure of spatiotemporal dynamics in the presence of symmetries through the use of invariant basis functions. The closure framework can be used to derive both deterministic and stochastic surrogate models; in this work we focus on the deterministic setting.
The main features of the present work can be summarized as follows.
  1. 1.
    Positivity preservation. As with the original QMCl work, our framework uses the noncommutative setting of operator algebras to encode the original classical dynamics, leading to a closure scheme that is positivity preserving. This means that positive flux terms in the original dynamics are represented by positive operators in our closure model, producing positive surrogate flux terms. Generating surrogate fluxes that preserve the sign of the original fluxes is a nontrivial task and has important consequences for the closure of dynamics with sign definite quantities.
  2. 2.
    Symmetry factorization. Taking advantage of the symmetry factorization properties of the VSA framework, our dynamical closure scheme rests on a representation of the dynamics that is invariant under the actions of the spatial symmetry group of the resolved dynamics. This enables the formation of a compressed representation of the original spatiotemporal dynamics, leading to a data driven realization of our closure scheme that can be applied to complex dynamics with tractable computational cost.
  3. 3.
    Computational efficiency. In addition to the factorization of spatial symmetries, we implement the data driven formulation of our scheme in a way that lowers the cost of its most computationally expensive operations. As with the original QMCl and VSA frameworks, our data driven scheme requires solving an eigenvalue problem for a kernel matrix. The dimension of this matrix can be very large, making the eigenvalue problem one of the most expensive tasks in our scheme. To address that we compute a low rank approximation of the original kernel matrix and use it to solve the eigenvalue problem with reduced cost [66, 67].
  4. 4.
    Application. We apply our scheme to a closure problem for the shallow water equations on a one dimensional spatial domain. The shallow water equations are a hyperbolic system of partial differential equations with important applications in fields such as water waves modeling and ocean dynamics [68, 69]. The results demonstrate that our scheme can effectively capture and reproduce the main features of the dynamics, including for out of sample initial conditions.
1.2 Paper layout
In Section 2 we introduce the closure problem under study. This is followed by a brief overview of the original QMCl framework in Section 3. In Section 4 we extend this framework to the closure of spatiotemporal dynamics. We present the main features of the extended scheme, the elements it borrows from the VSA framework, as well as its data driven realization. In Section 5 we apply the extended scheme to a closure problem for the shallow water equations, present our numerical results and discuss some aspects of the closure scheme and future work directions. This is followed by a brief conclusion in Section 6. Finally, A provides additional details on the numerical implementation of the data driven formulation of our closure scheme.
2 Problem statement
We consider a dynamical system $\Phi^{t}\colon X\rightarrow X$, $t\geq 0$ on a generally infinite dimensional state space $X$ possessing an invariant probability measure $\mu$. The flow map $\Phi^{t}$ represents the spatiotemporal dynamics governed by a partial differential equation on a bounded spatial domain $S\subset\mathbb{R}^{c}$ for some $c\in\mathbb{N}$. As such, elements of state space $X$ are functions on the domain $S$; more specifically, we consider $X\subset L^{2}(S,\nu;\mathbb{R}^{d})$ with $d\in\mathbb{N}$ and $\nu$ denoting the Lebesgue measure on $S$.
We decompose the state space as $X=\mathcal{X}\times\mathcal{Y}$ with $\mathcal{X}$ denoting the state space of the resolved degrees of freedom of the dynamics and $\mathcal{Y}$ the state space of the unresolved degrees. We assume that the resolved state space $\mathcal{X}\subset X$ is finite dimensional, a property that arises naturally in many applications. For instance many dissipative partial differential equations give rise to long time dynamics that is confined to a finite dimensional subset of their state space, such as an inertial manifold [70, 71]. In other cases discretization of the differential operator modeling the considered dynamics also yields a finite dimensional state space.
We define the Hilbert space of vector valued observables $H=L^{2}(X,\mu;H_{S})$ with $H_{S}=L^{2}(S,\nu_{M};\mathbb{C})$ where $\nu_{M}$ denotes a restriction of the Lebesgue measure $\nu$ that will be constructed in Section 4.1. Every observable $\vec{f}\in H$ represents a spatiotemporal pattern of the dynamics, for which every state $x\in X$ yields a function on the spatial domain $\vec{f}(x)\in H_{S}$.
We introduce $d$ real observables $z_{j}\in H$, $j\in\{0,\ldots,d-1\}$ which will allow us to include in the resolved dynamics terms that depend on the unresolved variables. We refer to the observables $z_{j}$ as the flux functions and to their generated terms $z_{j}(x)\in H_{S}$ for any $x\in X$ as their fluxes. In what follows we use $z=\{z_{j}\}_{j=0}^{d-1}$ to denote the collection of flux functions. Based on our decomposition of state space $X$ we define the projected flow maps for the resolved dynamics $\Phi_{\mathcal{X}}^{t}\colon X\to\mathcal{X}$, $\Phi_{\mathcal{X}}^{t}=\operatorname{proj}_{\mathcal{X}}\circ\Phi^{t}$ and unresolved dynamics $\Phi_{\mathcal{Y}}^{t}\colon X\to\mathcal{Y}$, $\Phi_{\mathcal{Y}}^{t}=\operatorname{proj}_{\mathcal{Y}}\circ\Phi^{t}$ where $\operatorname{proj}_{A}$ denotes the canonical projection onto a subspace $A\subset X$. With these definitions in place, the dynamics $\Phi^{t}$ can be expressed as
$\Phi^{t}(\hat{x},y)=\bigl{(}\Phi^{t}_{\mathcal{X}}(\hat{x},z(\hat{x},y)),\Phi_{\mathcal{Y}}^{t}(\hat{x},y)\bigr{)}$
with state vector $(\hat{x},y)\in\mathcal{X}\times\mathcal{Y}$.
Our goal in this work is to construct surrogate dynamics $\tilde{\Phi}^{t}\colon\tilde{X}\rightarrow\tilde{X}$ in a surrogate state space $\tilde{X}=\mathcal{X}\times\tilde{\mathcal{Y}}$ where $\tilde{\mathcal{Y}}$ denotes a surrogate unresolved state space. We introduce a surrogate version of the flux functions $\tilde{z}_{j}\colon\tilde{\mathcal{Y}}\to H_{S}$ and denote their collection by $\tilde{z}$. The surrogate dynamics $\tilde{\Phi}^{t}$ can now be expressed as
$\tilde{\Phi}^{t}(\hat{x},\tilde{y})=\bigl{(}\tilde{\Phi}_{\mathcal{X}}^{t}(\hat{x},\tilde{z}(\tilde{y})),\tilde{\Phi}_{\mathcal{Y}}^{t}(\hat{x},\tilde{y})\bigr{)}$ (1)
with state vector $(\hat{x},\tilde{y})\in\mathcal{X}\times\tilde{\mathcal{Y}}$, surrogate resolved dynamics $\tilde{\Phi}_{\mathcal{X}}^{t}\colon\tilde{X}\to\mathcal{X}$, $\tilde{\Phi}_{\mathcal{X}}^{t}=\operatorname{proj}_{\mathcal{X}}\circ\tilde{\Phi}$ and surrogate unresolved dynamics $\tilde{\Phi}_{\mathcal{Y}}^{t}\colon\tilde{X}\to\tilde{\mathcal{Y}}$, $\tilde{\Phi}_{\mathcal{Y}}^{t}=\operatorname{proj}_{\tilde{\mathcal{Y}}}\circ\tilde{\Phi}$. To build the surrogate dynamics map $\tilde{\Phi}^{t}$ we need to define the surrogate space $\tilde{\mathcal{Y}}$, the surrogate flux functions $\tilde{z}_{j}$ and the surrogate dynamics $\tilde{\Phi}_{\mathcal{X}}^{t}$ and $\tilde{\Phi}_{\mathcal{Y}}^{t}$.
We are particularly interested in systems with dynamical symmetries. For this reason we assume that there exists a symmetry group $G$ acting on the spatial domain $S$ with a continuous nontrivial left action $\Gamma_{S,g}\colon S\rightarrow S$, $g\in G$ which preserves the null sets of measure $\nu$. The actions $\Gamma_{S,g}$ induce the actions on state space $\Gamma_{X,g}\colon X\rightarrow X$, $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$, $g\in G$. The actions of group $G$ on $X$ represent dynamical symmetries of the dynamics $\Phi^{t}$; namely, we assume the following equivariance property
$\Gamma_{X,g}\circ\Phi^{t}=\Phi^{t}\circ\Gamma_{X,g}$
for all $t\geq 0$ and $g\in G$. In such situations we seek to leverage the dynamical symmetries to reduce training data requirements and the overall dimension of our surrogate models.
3 Quantum mechanical closure framework
In this section we provide a brief overview of QMCl as it has been developed in [56, 57, 55], which is concerned with the closure of dynamics governed by ordinary differential equations. The core of the QMCl framework consists of two steps: (1) embedding the classical dynamics into an infinite dimensional quantum mechanical system; (2) projecting the infinite dimensional system to a finite dimensional subsystem. The infinite dimensional embedding of the classical dynamics is treated in Section 3.1. In Section 3.2 we present the process used to project the embedded dynamics to a finite dimensional subsystem and derive a closure model.
In the description of QMCl provided below we omit some of the details required for its precise description and implementation. A detailed treatment of the scheme, including its data driven realization, will be given in the following sections once we extend it to the closure of dynamics governed by partial differential equations. A detailed analysis of QMCl as presented below can be found in [56, 57, 55].
3.1 Infinite dimensional embedding
To build the quantum mechanical closure framework we start by embedding the classical dynamics on state space $X$ into a quantum mechanical system on the infinite dimensional Hilbert space $H=L^{2}(X,\mu;\mathbb{C})$. This involves embedding observables (functions of state) and probability densities, as well as the evolution of said densities under the dynamics and their bayesian conditioning.
In the following we denote by $B(H)$ the space of bounded linear operators on Hilbert space $H$. We write $Q(H)$ to denote the space of quantum states on $H$ which are positive, trace class operators of unit trace. We use the term positive operator to describe a selfadjoint operator with nonnegative spectrum. Quantum states are also referred to as quantum density operators, and can be thought of as playing an analogous role to probability densities in classical statistics. In this work we mainly use pure quantum states, which are rank-1 quantum density operators and are equivalent to projections along a unit vector [58, 59, 60].
Every classical observable $f\in L^{\infty}(X,\mu;\mathbb{R})\subset H$ is mapped to a selfadjoint multiplication operator $A_{f}\in B(H)$, $A_{f}g=fg$ for all $g\in H$ referred to as a quantum observable. Additionally, every classical probability density $p\in L^{1}(X,\mu;\mathbb{R})$, $\sqrt{p}\in H$ is mapped to the quantum density operator $\rho_{p}\in Q(H)$, $\rho_{p}=\langle\sqrt{p},\cdot\rangle\sqrt{p}$ representing a pure quantum state. The expected value of observable $A\in B(H)$ when the quantum system is in state $\rho\in Q(H)$ is given by $\mathbb{E}_{\rho}A=\operatorname{tr}(\rho A)\in\mathbb{R}$ which is connected to the theory of quantum measurement [58, 59, 60].
We define the surrogate unresolved state space $\tilde{\mathcal{Y}}=Q(H)$ and use a quantum state $\rho\in Q(H)$ to model the state of the unresolved degrees of freedom of the dynamics. In addition, we define the operators $A_{j}\in B(H)$, $j\in\{0,\ldots,d-1\}$ to model $d$ observables of the system’s unresolved state, which are constructed by mapping the true flux functions $z_{j}\in H$ to multiplication operators. With these definitions in place we introduce the surrogate flux functions $\tilde{z}_{j}\colon Q(H)\rightarrow\mathbb{R}$, $j\in\{0,\ldots,d-1\}$ with action
$\tilde{z}_{j}(\rho)=\mathbb{E}_{\rho}(A_{j})=\operatorname{tr}(\rho A_{j}).$
In other words we use the space of quantum states $\tilde{\mathcal{Y}}=Q(H)$ to model the unresolved degrees of freedom of the dynamics and the framework of quantum measurement to compute their fluxes.
In addition to evaluating the required flux terms we need to be able to evolve a quantum state under the system’s dynamics and condition it based on observed dynamical behavior. In the classical case, the square root of probability densities $\sqrt{p}\in H$ evolves according to the unitary group of transfer operators $P^{t}\colon H\to H$, $P^{t}\sqrt{p}=\sqrt{p}\circ\Phi^{-t}$ with $t\geq 0$. Accordingly, the evolution of quantum states $\rho\in Q(H)$ is governed by the induced transfer operator $\mathcal{P}^{t}\colon Q(H)\to Q(H)$, $\mathcal{P}^{t}\rho=P^{t}\rho P^{-t}$. In analogy to bayesian conditioning of classical probability densities, we can condition a quantum state $\rho\in Q(H)$ using a quantum effect operator $e\in B(H)$, $0\leq e\leq\operatorname{Id}$
$\rho|_{e}=\frac{\sqrt{e}\rho\sqrt{e}}{\mathbb{E}_{\rho}e}$ (2)
where $\sqrt{e}$ denotes the positive square root of $e$. The above equation can be interpreted as the quantum analogue of the classical Bayes rule [55].
A quantum effect operator can be thought of as the quantum analogue of an indicator function used in classical probability to represent an event, namely a set in the event space. For instance, for a given set $F\subset X$ we can build the indicator function $1_{F}\in H$ taking the value one on $F$ and zero elsewhere. The quantum effect operator $e_{F}\in B(H)$ can then be defined as multiplication by $1_{F}$, namely $e_{F}f=1_{F}f$ for all $f\in H$.
3.2 Closure modeling
We build a finite dimensional subspace $H^{L}\subset H$ using the leading $L$ elements of an orthonormal basis $\{\phi_{\ell}\}_{\ell=0}^{\infty}$ of $H$. To obtain such a basis we define a positive kernel integral operator $K\colon H\to H$
[…]
using a positive, symmetric and uniformly continuous kernel function $\kappa\in C_{b}(X\times X;\mathbb{R})\cap L^{2}(X\times X,\mu\times\mu;\mathbb{R})$ where $C_{b}$ denotes the space of bounded continuous functions. Operator $K$ is compact and selfadjoint, so its eigenfunctions form an orthonormal basis of $H$; we write $K\phi_{\ell}=\lambda_{\ell}\phi_{\ell}$ where the eigenvalues $\lambda_{\ell}$ are nonnegative and ordered in decreasing order. We define $H^{L}$ as the linear span of the leading $L$ eigenfunctions of operator $K$. Having built $H^{L}$ we project operators acting on $H$ to operators acting on $H^{L}$; that is, $A\in B(H)$ is approximated by $A_{L}=\operatorname{proj}_{L}A\operatorname{proj}_{L}$ where $\operatorname{proj}_{L}\in B(H)$ is the orthogonal projection onto $H^{L}$. This is done for all operators needed for the quantum mechanical embedding of the original classical dynamics.
As explained earlier, to build our closure scheme we set $\tilde{\mathcal{Y}}=Q(H^{L})$ and use quantum states to model the unresolved degrees of freedom of the surrogate dynamics $\tilde{\Phi}^{t}$ from (1). We now describe the calculation of the fluxes needed for evolving the resolved state under $\tilde{\Phi}_{\mathcal{X}}^{t}$, and the surrogate map $\tilde{\Phi}_{\mathcal{Y}}^{t}$ used to evolve the quantum state representing the unresolved state. To do so, we use a discrete time version of the above flow maps, denoting them by $\tilde{\Phi}_{\mathcal{X}}^{n\Delta t}$ and $\tilde{\Phi}_{\mathcal{Y}}^{n\Delta t}$ with $n\in\mathbb{N}$ and timestep $\Delta t>0$.
Given the resolved system state $\hat{x}_{n}\in\mathcal{X}$ at time $t_{n}=n\Delta t$ we use a quantum state $\rho_{n}\in Q(H^{L})$ to represent the unresolved state and $d$ quantum observables $A_{j}\in B(H^{L})$, $j\in\{0,\ldots,d-1\}$. Each observable $A_{j}$ is a projected multiplication operator defined using the corresponding true flux function $z_{j}\in H$. We seek to compute state $\hat{x}_{n+1}=\tilde{\Phi}_{\mathcal{X}}(\hat{x}_{n},\tilde{z}(\rho_{n}))$ using the collection of surrogate flux functions $\tilde{z}=\{\tilde{z}_{j}\}_{j=0}^{d-1}$ with $\tilde{z}_{j}\colon Q(H^{L})\to\mathbb{R}$. Each surrogate flux term $\tilde{z}_{j}(\rho_{n})$ is computed as the expectation value of observable $A_{j}$ with respect to quantum state $\rho_{n}$
$\tilde{z}_{j}(\rho_{n})=\mathbb{E}_{\rho_{n}}(A_{j})=\operatorname{tr}{(\rho_{n}A_{j})}.$
The resolved state is then updated by $\hat{x}_{n+1}=\tilde{\Phi}_{\mathcal{X}}^{\Delta t}(\hat{x}_{n},\tilde{z}(\rho_{n}))$.
Next we define the surrogate unresolved dynamics map $\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}$ used to update the quantum state $\rho_{n+1}=\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}(\hat{x}_{n},\rho_{n})$. The quantum state update consists of two steps, a prediction and a correction step. Given the current quantum state $\rho_{n}$ we first use the projected transfer operator $\mathcal{P}^{L}\colon Q(H^{L})\rightarrow Q(H^{L})$ to compute the prior state
$\rho_{n+1}^{\prime}=\frac{\mathcal{P}^{L}\rho_{n}}{\operatorname{tr}{(\mathcal{P}^{L}\rho_{n})}}.$
Renormalizing the quantum state after acting with $\mathcal{P}^{L}$ is required because the projected operator $\mathcal{P}^{L}$ is generally not unitary. The mapping $\rho_{n}\mapsto\rho_{n+1}^{\prime}$ corresponds to the prediction step. The correction step involves using the updated resolved state $\hat{x}_{n+1}$ to update our prediction $\rho_{n+1}^{\prime}$. To do so, we use state $\hat{x}_{n+1}$ to build a quantum effect operator $e_{n+1}\in B(H^{L})$; we refer to the mapping $\hat{x}_{n+1}\mapsto e_{n+1}$ as the feature map. Using the quantum Bayes rule (2) we condition the prior state $\rho_{n+1}^{\prime}$ to compute the posterior state
$\rho_{n+1}=\rho_{n+1}^{\prime}|_{e_{n+1}}=\frac{\sqrt{e_{n+1}}\rho_{n+1}^{\prime}\sqrt{e_{n+1}}}{\operatorname{tr}{(\sqrt{e_{n+1}}\rho_{n+1}^{\prime}\sqrt{e_{n+1}}})}.$
The above two steps make up the surrogate unresolved dynamics $\tilde{\Phi}_{\mathcal{Y}}^{\Delta t}$. Although not unitary, operator $\mathcal{P}^{L}$ is trace nonincreasing which means that it can be interpreted as governing the evolution of an open quantum system $H^{L}\subset H$. Under this light, the reduction of $\operatorname{tr}(\rho)$ when acting with $\mathcal{P}^{L}$ can be viewed as decoherence or leakage of density due to interaction with the environment [58, 59, 60].
The QMCl closure framework as developed above can be summarized in the following diagram, which depicts the evolution of the built surrogate dynamical system for one timestep [55].
[…]
In the diagram $\operatorname{Id}$ denotes the identity map.
4 Closure of partial differential equations
In this section we extend the QMCl scheme to the case of spatiotemporal dynamics governed by partial differential equations. One of the distinguishing features of the extended scheme is the appropriate handling of dynamical symmetries. We develop the extended scheme by pairing QMCl with the vector valued spectral analysis (VSA) framework [65], which combines methods from operator valued kernels and delay embedding to construct efficient compressed representations of dynamics in the presence of dynamical symmetries.
In Section 4.1 we define the appropriate Hilbert space $H$ and describe the process of embedding the classical spatiotemporal dynamics into a quantum mechanical system in $H$. In Section 4.2 we introduce the kernel used to discretize (compress) the embedded dynamics to a finite dimensional subsystem. Section 4.3 explores the symmetry invariance property of the derived eigenfunction basis. Finally, Sections 4.4 and 4.5 present the data driven formulation of the extended QMCl scheme that can be realized numerically.
4.1 Closure modeling
As explained in Section 2 we consider spatiotemporal dynamics $\Phi^{t}\colon X\rightarrow X$ possessing an invariant probability measure $\mu$ and a generally infinite dimensional state space $X\subset L^{2}(S,\nu;\mathbb{R}^{d})$ where $S$ denotes a bounded spatial domain. We assume the state space decomposition $X=\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}$ is the finite dimensional resolved state space and $\mathcal{Y}$ the unresolved state space.
To provide a general definition of the resolved state space $\mathcal{X}\subset X$ we consider the measure space $(S,\mathcal{A},\nu)$ where $\mathcal{A}$ denotes the Borel $\sigma$-algebra on $S$. We form the subalgebra $\mathcal{A}_{M}\subset\mathcal{A}$ generated by a finite collection of $M\in\mathbb{N}$ disjoint subsets $S_{m}\subset S$, $m\in\{0,\ldots,M-1\}$ such that $\cup_{m=0}^{M-1}S_{m}=S$. Next we define measure $\nu_{M}$ as the restriction of $\nu$ to $\sigma$-algebra $\mathcal{A}_{M}$. The resolved state space is defined as the $dM$-dimensional subspace $\mathcal{X}=L^{2}(S,\nu_{M};\mathbb{R}^{d})\cap X$. As a Hilbert space, $\mathcal{X}$ is isomorphic to $\mathbb{R}^{dM}$ with a scaled euclidean inner product. Each subset $S_{m}\subset S$ used to generate subalgebra $\mathcal{A}_{M}$ can also be represented by its center point $s_{m}\in S_{m}$. We denote by $S_{M}=\{s_{m}\}_{m=0}^{M-1}\subset S$ the collection of those center points. Every equivalence class of functions $\tilde{h}\in L^{2}(S,\nu_{M};\mathbb{R}^{d})$ has a piecewise constant representative $h=\sum_{j=0}^{M-1}\alpha_{j}1_{S_{j}}$ where $1_{S_{j}}$ denotes the indicator function of set $S_{j}\in\{S_{m}\}_{m=0}^{M-1}$ and $\alpha_{j}\in\mathbb{R}^{d}$ the corresponding coefficient.
As our quantum mechanical Hilbert space $H$ we consider the space of vector valued observables $H=L^{2}(X,\mu;H_{S})$ with $H_{S}=L^{2}(S,\nu_{M};\mathbb{C})$. Introducing the product state space $\Omega=X\times S$ and measure $\sigma=\mu\times\nu_{M}$ we define $H_{\Omega}=L^{2}(\Omega,\sigma;\mathbb{C})$ and note the following Hilbert space isomorphisms
$H\simeq H_{X}\otimes H_{S}\simeq H_{\Omega},$
where $H_{X}=L^{2}(X,\mu;\mathbb{C})$. Namely, for every vector valued observable $\vec{f}\in H$ there is a corresponding complex valued $f\in H_{\Omega}$ such that for almost every $x\in X$ and $s\in S$ $f(x,\cdot)=\vec{f}(x)\in H_{S}$ and $f(x,s)=\vec{f}(x)(s)\in\mathbb{C}$. The motivation for our choice of space $H$ is our interest in observables that can effectively capture spatiotemporal dynamics; that is, patterns of evolution that have both a temporal dependence $x\in X$ and a spatial dependence $s\in S$. Given an observable $\vec{f}\in H$ the map $t\mapsto\vec{f}(\Phi^{t}(x))$ represents a spatiotemporal pattern generated by the dynamics $\Phi^{t}$ [65].
The definitions of quantum observables, quantum effects and the induced transfer operator carry over from Section 3 after using the new definition of Hilbert space $H$. To model the unresolved degrees of freedom we introduce the surrogate unresolved state space $\tilde{\mathcal{Y}}=L^{2}(S,\nu_{M};B(H))$. Namely, we use a field of quantum states over $S$, $\rho\in\tilde{\mathcal{Y}}$ with $\operatorname{ran}{\rho}\subseteq Q(H)\subset B(H)$, where each quantum state $\rho(s)$ represents the unresolved degrees of freedom of the dynamics at point $s\in S$ of the spatial domain. This allows us to compute fluxes that vary over the spatial domain $S$ for each quantum observable $A_{j}\in B(H)$, $j\in\{0,\ldots,d-1\}$. More specifically, we define the corresponding $d$ surrogate flux functions $\tilde{z}_{j}\colon\tilde{\mathcal{Y}}\to H_{S}$ with action
$\tilde{z}_{j}(\rho)(s)=\operatorname{tr}(\rho(s)A_{j})\in\mathbb{R}$
yielding $d$ flux terms at each spatial point.
Having embedded the original dynamics into an infinite dimensional quantum system in Hilbert space $H$, we now have to discretize space $H$ using the eigendecomposition of a kernel integral operator. We tackle this task in the next section.
4.2 Choice of kernel
To build our closure scheme on $H$ we use data generated from a dynamical trajectory $x_{n}=\Phi^{n\Delta t}(x_{0})$ with $x_{0}\in X$, $n\in\mathbb{N}$ and sampling timestep $\Delta t>0$. We do not require access to the full state samples $x_{n}\in X$; only to the corresponding resolved state samples $\hat{x}_{n}=\operatorname{proj}_{\mathcal{X}}(x_{n})\in\mathcal{X}$ and flux samples $z_{j}(x_{n})\in H_{S}$, $j\in\{0,\ldots,d-1\}$. Note that for each $n\in\mathbb{N}$ $\hat{x}_{n}$ and $z_{j}(x_{n})$ are functions on $S$ with $\hat{x}_{n}(s)$ and $z_{j}(x_{n})(s)$ denoting their values at point $s\in S$ of the spatial domain.
To discretize the space $H_{\Omega}\simeq H$ we use the orthonormal basis generated by the eigenfunctions of a positive kernel integral operator $K\colon H_{\Omega}\rightarrow H_{\Omega}$
[…]
with a positive, symmetric and uniformly continuous kernel $\kappa\in C_{b}(\Omega\times\Omega;\mathbb{R})\cap L^{2}(\Omega\times\Omega,\sigma\times\sigma;\mathbb{R})$. By forming an operator acting directly on $H_{\Omega}$ we obtain eigenfunctions that are generally not of separable, tensor product form. This is in contrast to traditional approaches such as the proper orthogonal decomposition where one builds a basis for $H_{X}$ (the temporal space in this context) and forms its tensor product with a basis for the spatial space $H_{S}$ to form a basis for $H_{\Omega}\simeq H_{X}\otimes H_{S}$ [72]. As we will see below, in the presence of dynamical symmetries our approach leads to a more compressed finite dimensional representation of the dynamics on $H$, meaning that each eigenfunction of operator $K$ as defined above can represent spatiotemporal patterns that are not expressible as a product of a single pair of temporal and spatial modes.
To have a kernel function $\kappa$ that relies only on the sampled information we have access to we employ data space $\mathcal{W}$ and data map $W\colon\Omega\to\mathcal{W}$. We introduce the data space kernel $k\colon\mathcal{W}\times\mathcal{W}\to\mathbb{R}$ and consider $\kappa$ as its pullback $\kappa(\omega,\omega^{\prime})=k(W(\omega),W(\omega^{\prime}))$. In this work we use a kernel $k$ that acts in delay coordinate space $\mathcal{W}=\mathbb{R}^{Q}$ where $Q\in\mathbb{N}$ denotes the number of delays. As explained below, for each sample $\omega=(x,s)\in X\times S$ we embed the resolved state component $(\operatorname{proj}_{\mathcal{X}}(x),s)\in\mathcal{X}\times S$ into $\mathbb{R}^{Q}$ by using delays that act locally at spatial point $s\in S$. This allows us to recover some of the dynamical information lost when projecting to the resolved component of the state as well as to factor redundancies introduced in the sampled data by dynamical symmetries [65].
Consider the data map $W_{Q}\colon\Omega\rightarrow\mathbb{R}^{Q}$
$W_{Q}((x,s))=\bigl{(}\operatorname{proj}_{\mathcal{X}}(x)(s),\,\operatorname{proj}_{\mathcal{X}}(\Phi^{-\Delta t}(x))(s),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(\Phi^{-(Q-1)\Delta t}(x))(s)\bigr{)}$ (4)
which corresponds to $S$-pointwise delay embedding of product state sample $\omega=(x,s)$. Using this map we define the gaussian kernel function $k\colon\mathbb{R}^{Q}\times\mathbb{R}^{Q}\to\mathbb{R}$
$k(W_{Q}(\omega),W_{Q}(\omega^{\prime}))=\exp\Bigl{(}-\frac{1}{\epsilon Q}\lVert W_{Q}(\omega)-W_{Q}(\omega^{\prime})\rVert_{Q}^{2}\Bigr{)}$ (5)
where $\lVert\cdot\rVert_{Q}$ denotes the 2-norm in $\mathbb{R}^{Q}$ and $\epsilon>0$ is a tunable bandwidth parameter. The parameter $\epsilon$ can be calibrated using the procedure developed in [73]. In our numerical experiments we normalize the above kernel to turn it into a bistochastic kernel. The bistochastic kernel normalization process and motivation for using it can be found in [74, 75].
With this choice of kernel function we use the eigendecomposition of operator $K$ from (3) to build an orthonormal basis for its range $\operatorname{ran}{K}\subset H_{\Omega}$. We then use its $L$ leading eigenfunctions $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ to define the finite dimensional subspace $H_{\Omega}^{L}$ as their linear span. Once subspace $H_{\Omega}^{L}$ has been formed we proceed to project all operators involved in QMCl from $B(H_{\Omega})$ to $B(H_{\Omega}^{L})$. The resulting operators in $B(H_{\Omega}^{L})$ can be represented by $L\times L$ matrices with respect to our basis. This process will be developed in Section 4.4 along with the data driven realization of the extended closure scheme.
In the next section we describe the properties of the subspace $\operatorname{ran}{K}\subset H_{\Omega}$ obtained through the eigendecomposition of $K$.
4.3 Dynamical symmetries
We consider a symmetry group $G$ acting on the spatial domain $S$ by a continuous left action $\Gamma_{S,g}\colon S\rightarrow S$, $g\in G$. As explained in Section 2 we assume that the dynamics $\Phi^{t}$ is equivariant under the actions of $G$
$\Gamma_{X,g}\circ\Phi^{t}=\Phi^{t}\circ\Gamma_{X,g}$
for all $t\geq 0$ and $g\in G$, where $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$ denotes the group’s induced action on state space $X$. Every action $\Gamma_{X,g}$ represents a dynamical symmetry of the dynamics $\Phi^{t}$.
We consider the decomposition of dynamics $\Phi^{t}$ into its resolved $\Phi_{\mathcal{X}}^{t}$ and unresolved $\Phi_{\mathcal{Y}}^{t}$ components. In general, the coarsening process used to derive the resolved dynamics from the original may lead to loss of some of the original symmetries. Accordingly, we focus on a subgroup $\mathcal{G}\subseteq G$ that corresponds to the set of symmetries obeyed by the resolved dynamics $\Phi_{\mathcal{X}}^{t}$. We define the induced action $\Gamma_{\mathcal{X},g}\colon\mathcal{X}\rightarrow\mathcal{X}$, $\Gamma_{\mathcal{X},g}=\operatorname{proj}_{\mathcal{X}}\circ\Gamma_{X,g}$, $g\in\mathcal{G}$ and assume that the resolved dynamics $\Phi_{\mathcal{X}}^{t}$ is equivariant under the appropriate actions of $\mathcal{G}$
$\Gamma_{\mathcal{X},g}\circ\Phi_{\mathcal{X}}^{t}=\Phi_{\mathcal{X}}^{t}\circ\Gamma_{X,g}$ (6)
for all $t\geq 0$ and $g\in\mathcal{G}$. In addition, we assume that the resolved dynamics is derived in such a way that the true flux functions $z_{j}\in H$ satisfy the analogous equivariance property
$\Gamma_{z,g}\circ z_{j}=z_{j}\circ\Gamma_{X,g}$
for all $g\in\mathcal{G}$ and $j\in\{0,\ldots,d-1\}$ with induced action $\Gamma_{z,g}\colon H_{S}\to H_{S}$. From now on we focus on subgroup $\mathcal{G}$ and equivariance property (6). As explained below, our choice of kernel naturally exploits this equivariance property to bestow an analogous invariance property on the eigenfunctions of the kernel integral operator.
Under the action of data map $W_{Q}$ from (4) the product space $\Omega$ can be viewed as a topological fiber bundle $(\Omega,B_{Q},W_{Q})$, where $\Omega$ is the total space, $B_{Q}=W_{Q}(\Omega)$ the base space and $W_{Q}\colon\Omega\rightarrow B_{Q}$ the continuous projection map. The fibers are equivalence classes $[\omega]_{Q}=W_{Q}^{-1}(b)\subseteq\Omega$, $b\in B_{Q}$ defined as the set of values on which $W_{Q}$ is constant. For our choice of projection map $W_{Q}$ each fiber consists of points in $\Omega$ whose resolved dynamical evolution over $Q$ delays is identical.
Using the group action $\Gamma_{\Omega,g}\colon\Omega\rightarrow\Omega$, $\Gamma_{\Omega,g}=\Gamma_{X,g}\otimes\Gamma_{S,g}$ we define the $\mathcal{G}$-orbit of $\omega\in\Omega$ as the set $\Gamma_{\Omega}(\omega)=\{\Gamma_{\Omega,g}(\omega)\colon g\in\mathcal{G}\}$. As shown in [65], $\Gamma_{\Omega}(\omega)\subseteq[\omega]_{Q}$ for all $\omega\in\Omega$; that is, every fiber $[\omega]_{Q}$ contains the symmetric transformations of $\omega$ induced by group $\mathcal{G}$.
Because kernel function $\kappa$ in (3) acts by composition with the projection map $W_{Q}$, $\kappa=k\circ(W_{Q}\times W_{Q})$, it is constant on the $[\cdot]_{Q}$ fibers by construction. As a result, every function $f\in\operatorname{ran}{K}$ is constant on the fibers; namely, $f(\omega^{\prime})=f(\omega)$ for all $\omega^{\prime}\in[\omega]_{Q}$. Given that the $\mathcal{G}$-orbits $\Gamma_{\Omega}(\omega)\subseteq[\omega]_{Q}$ for all $\omega\in\Omega$, it follows that functions $f\in\operatorname{ran}{K}$ satisfy the invariance property
$f\circ\Gamma_{\Omega,g}=f$ (7)
for all $g\in\mathcal{G}$. In particular this implies that all eigenfunctions of integral operator $K$ corresponding to nonzero eigenvalues satisfy this property as well. Combined with the fact that all $f\in\operatorname{ran}{K}$ are continuous by the uniform continuity of $\kappa$, we conclude that $\operatorname{ran}{K}$ consists entirely of continuous functions that are constant on the fibers $[\cdot]_{Q}$.
As a result of (7), by using the eigenfunctions of $K$ corresponding to nonzero eigenvalues we derive a basis of $\operatorname{ran}{K}$ that is invariant under the actions of $\mathcal{G}$, instead of getting multiple “copies” of the same underlying functions generated by the action of different $g\in\mathcal{G}$. Most importantly for our purposes, this means that fewer eigenfunctions are needed to represent complex spatiotemporal dynamics, thereby producing a more efficient compression of the dynamics from $H_{\Omega}$ to subspace $H_{\Omega}^{L}$ [65].
4.4 Data driven formulation
In this section we use the framework developed in the previous sections to define a data driven closure scheme that can be realized numerically. We first establish the data driven formulation using one dynamical trajectory to generate our training data. In the next section we extend this to the case of training data consisting of multiple trajectories.
We consider a finite dynamical trajectory of $N\in\mathbb{N}$ samples $X_{N}=\{x_{n}\}_{n=0}^{N-1}\subset X$ with $x_{n}=\Phi^{n\Delta t}(x_{0})$, $x_{0}\in X$, $\Delta t>0$. We define the discrete sampling measure $\mu_{N}=\sum_{n=0}^{N-1}\delta_{x_{n}}/N$ and corresponding finite dimensional Hilbert space $H_{X,N}=L^{2}(X,\mu_{N};\mathbb{C})$ where functions are represented by their values on samples $X_{N}\subset X$. Combining the above we define the associated product space $H_{\Omega,N}=L^{2}(\Omega,\sigma_{N};\mathbb{C})$ with discrete product measure $\sigma_{N}=\mu_{N}\times\nu_{M}$.
[…]
As before we do not assume access to the dynamical trajectory samples $x_{n}$ but only to the resolved state samples $\hat{x}_{n}=\operatorname{proj}_{\mathcal{X}}(x_{n})\in\mathcal{X}$ and flux samples $z_{j}(x_{n})\in H_{S}$, $j\in\{0,\ldots,d-1\}$. We denote the state samples in product form by $\omega_{nm}=(x_{n},s_{m})\in\Omega$ with state $x_{n}\in X_{N}$ and gridpoint $s_{m}\in S_{M}$.
Using the data space kernel $k$ of (5) we define the associated data driven kernel integral operator $K_{N}\colon H_{\Omega,N}\rightarrow H_{\Omega,N}$
[…]
Using the $L$ leading eigenfunctions $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ of $K_{N}$ we construct the $L$-dimensional subspace $H_{\Omega,N}^{L}\subset H_{\Omega,N}$ as their linear span. Computing the $L$ leading eigenfunctions of $K_{N}$ can become very expensive as the dimension of the product space $NM$ is increased. To address that, in our numerical experiments we build a low rank approximation of the kernel matrix $\bm{K}_{N}\in\mathbb{R}^{NM\times NM}$ associated with operator $K_{N}$ to compute its eigenfunctions with reduced cost (Section 5.3). In what follows we make use of the orthogonal projection $\operatorname{proj}_{L}\colon H_{\Omega,N}\rightarrow H_{\Omega,N}$, $\operatorname{ran}{\operatorname{proj}_{L}}=H_{\Omega,N}^{L}$. Using $\operatorname{proj}_{L}$ an operator $A\in B(H_{\Omega,N})$ can be projected by conjugation $\operatorname{Proj}_{L}A=\operatorname{proj}_{L}A\operatorname{proj}_{L}\in B(H_{\Omega,N}^{L})$. Note that the latter projection preserves positivity; namely, $\operatorname{Proj}_{L}A$ is a positive operator whenever $A$ is positive.
In the data driven scheme quantum observables are selfadjoint operators in $B(H_{\Omega,N}^{L})$. To construct such an observable $A_{j}$ we first define a multiplication operator $M_{f}\in B(H_{\Omega,N})$, $M_{f}g=fg$ for all $g\in H_{\Omega,N}$ where $f(\omega_{nm})=z_{j}(x_{n})(s_{m})$ corresponds to the true flux function $z_{j}$ viewed as a real valued function in $H_{\Omega,N}$. Orthogonal projection then yields $A_{j}=\operatorname{Proj}_{L}M_{f}\in B(H_{\Omega,N}^{L})$. Note that $A_{j}$ is generally no longer a multiplication operator. Using the orthonormal basis $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ each operator $A_{j}\in B(H_{\Omega,N}^{L})$ is represented by a matrix $\bm{A}_{j}\in\mathbb{R}^{L\times L}$ with entries $A_{j,ik}=\langle\phi_{i},A_{j}\phi_{k}\rangle_{N}$. In this way for any $g\in H_{\Omega,N}^{L}$ the operation $A_{j}g$ is represented by $\bm{A}_{j}\underline{g}$ with coefficient vector $\underline{g}\in\mathbb{R}^{L}$, $g_{\ell}=\langle\phi_{\ell},g\rangle_{N}$. Two distinguishing features of the discretization process employed by QMCL is that it is contractive and positivity preserving. The first property means that the surrogate fluxes produced by operator $A_{j}$ are bounded in absolute value by the maximum absolute value of $z_{j}\in H_{\Omega,N}$. The second property means that that if function $z_{j}$ is sign definite (positive or negative) then so will be operator $A_{j}$, producing surrogate fluxes that agree in sign with the true fluxes.
Quantum states form a field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$ with one quantum state $\rho_{m}=\rho(s_{m})\in Q(H_{\Omega,N}^{L})\subset B(H_{\Omega,N}^{L})$ for each gridpoint $s_{m}\in S_{M}$. In this work we focus on pure quantum states, namely rank-1 quantum density operators acting as projections along a given vector. For a gridpoint $s_{m}\in S_{M}$ quantum state $\rho_{m}$ is constructed as the projection along a real valued function $h_{m}\in H_{\Omega,N}^{L}$, $\rho_{m}=\langle h_{m},\cdot\rangle_{N}h_{m}$. As such, we represent each state $\rho_{m}$ by the coefficient vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$, $\rho_{m,\ell}=\langle\phi_{\ell},h_{m}\rangle_{N}$. This is equivalent to representing $\rho_{m}$ by the matrix $\bm{\rho}_{m}\in\mathbb{R}^{L\times L}$, $\rho_{m,ik}=\langle\phi_{i},\rho_{m}\phi_{k}\rangle_{N}$ but leads to savings in memory and computation.
For each $j\in\{0,\ldots,d-1\}$ the surrogate flux function $\tilde{z}_{j}\colon L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))\to H_{S}$ maps the quantum state field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L})))$ to the scalar field $\tilde{z}_{j}(\rho)\in H_{S}$ with $\tilde{z}_{j}(\rho)(s_{m})\in\mathbb{R}$ the computed flux at gridpoint $s_{m}\in S_{M}$. More specifically, given $d$ quantum observables $A_{j}\in B(H_{\Omega,N}^{L})$ and a quantum state field $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$ each surrogate flux function $\tilde{z}_{j}$ acts by $\tilde{z}_{j}(\rho)(s_{m})=\operatorname{tr}(\rho(s_{m})A_{j})\in\mathbb{R}$. Using our basis representation of each $A_{j}$ by matrix $\bm{A}_{j}\in\mathbb{R}^{L\times L}$ and $\rho(s_{m})$ by vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$ we have that $\tilde{z}_{j}(\rho)(s_{m})=\underline{\rho}_{m}^{\top}\bm{A}_{j}\underline{\rho}_{m}$.
The evolution of quantum states is governed by the projected induced transfer operator $\mathcal{P}_{N}^{L}\colon B(H_{\Omega,N}^{L})\to B(H_{\Omega,N}^{L})$. Because we work with pure quantum states—represented by vectors instead of matrices—we are going to form the projected transfer operator $P_{N}^{L}\colon H_{\Omega,N}^{L}\to H_{\Omega,N}^{L}$ instead of $\mathcal{P}_{N}^{L}$. To build $P_{N}^{L}$ we define the shift operator $P_{N}\colon H_{\Omega,N}\to H_{\Omega,N}$
$P_{N}f(\hat{x}_{n},s_{m})=\begin{cases}0&\text{if $n=0$}\\ f(\hat{x}_{n-1},s_{m})&\text{if $1\leq n<N$}.\end{cases}$
After projection we obtain $P_{N}^{L}=\operatorname{Proj}_{L}P_{N}$ represented by a matrix $\bm{P}_{N}^{L}\in\mathbb{R}^{L\times L}$, $(P_{N}^{L})_{ik}=\langle\phi_{i},P_{N}^{L}\phi_{k}\rangle_{N}$. Given a single quantum state $\rho_{n}\in Q(H_{\Omega,N}^{L})$ at time $t_{n}$ with vector representation $\underline{\rho}_{n}\in\mathbb{R}^{L}$ we obtain the updated state $\underline{\rho}_{n+1}$ by
$\underline{\rho}_{n+1}=\frac{\bm{P}_{N}^{L}\,\underline{\rho}_{n}}{\lVert\bm{P}_{N}^{L}\,\underline{\rho}_{n}\rVert_{2}}.$
The matrix $\bm{P}_{N}^{L}$ is in general not unitary, which is why we renormalize the updated quantum state after applying $\bm{P}_{N}^{L}$.
Next, quantum effects are operators $e\in B(H_{\Omega,N}^{L})$ used to condition quantum states based on a given state vector $\omega\in\Omega$. To build effect $e$ from state vector $\omega$ we introduce the conditioning kernel function $\kappa_{c}\colon\Omega\times\Omega\to\mathbb{R}$, defined as the pullback of kernel $k_{c}\colon\mathcal{W}_{c}\times\mathcal{W}_{c}\to\mathbb{R}$, $\kappa_{c}(\omega,\omega^{\prime})=k_{c}(W_{c}(\omega),W_{c}(\omega^{\prime}))$ with conditioning data space $\mathcal{W}_{c}$ and associated data map $W_{c}\colon\Omega\to\mathcal{W}_{c}$. To define kernel $k_{c}$ we introduce the data space $\mathcal{W}_{c}=\mathbb{R}^{J}$ with parameter $J\in\mathbb{N}$. Additionally, we define the neighborhood function $\mathcal{N}\colon S\to S^{J}$ which for a given gridpoint $s_{m}\in S_{M}$ returns its $J$ nearest neighbors on $S_{M}$ in a specified order. For a given gridpoint $s_{m}\in S_{M}$ and state vector $\omega_{m}=(x,s_{m})\in\Omega$ we then define the action of data map $W_{J}\colon\Omega\to\mathbb{R}^{J}$ by
$W_{J}((x,s_{m}))=\bigl{(}\operatorname{proj}_{\mathcal{X}}(x)(\mathcal{N}^{(0)}(s_{m})),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(x)(\mathcal{N}^{(J-1)}(s_{m}))\bigr{)}$
where $\mathcal{N}^{(j)}(s_{m})$ denotes the $j$th component of vector $\mathcal{N}(s_{m})$. For example, for a one dimensional grid $S_{M}$ and odd $J\in\mathbb{N}$ the action of $W_{J}$ becomes
$W_{J}((x,s_{m}))=\bigl{(}\operatorname{proj}_{\mathcal{X}}(x)(s_{m-(J-1)/2}),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(x)(s_{m}),\,\ldots,\,\operatorname{proj}_{\mathcal{X}}(x)(s_{m+(J-1)/2})\bigr{)}.$
The kernel $k_{c}\colon\mathbb{R}^{J}\times\mathbb{R}^{J}\to\mathbb{R}$ is then defined as
$k_{c}(W_{J}(\omega),W_{J}(\omega^{\prime}))=\exp\bigl{(}-\frac{1}{\epsilon^{\prime}J}\lVert W_{J}(\omega)-W_{J}(\omega^{\prime})\rVert_{J}^{2}\bigr{)}$ (9)
with $\lVert\cdot\rVert_{J}$ the 2-norm in $\mathbb{R}^{J}$ and $\epsilon^{\prime}>0$ a bandwidth parameter. Note that the bandwidth parameter $\epsilon^{\prime}$ is different from the one introduced earlier in kernel (5); it can again be selected using the tuning procedure developed in [73]. In our numerical experiments we employ the above kernel using a variable bandwidth function; details on the bandwidth function and an algorithm for computing it can be found in [76, 64].
For a fixed gridpoint $s_{m}\in S_{M}$ and given a state vector $\omega_{m}$ and the full training dataset $\{\omega_{nm}\}_{nm=0}^{NM-1}$, kernel $\kappa_{c}$ is used to produce a function $f_{m}\in H_{\Omega,N}$, $f_{m}(\cdot)=\kappa_{c}(\omega_{m},\cdot)$ leading to multiplication operator $M_{f_{m}}$ and the desired quantum effect $e_{m}=\operatorname{Proj}_{L}M_{f_{m}}$. Employing the quantum Bayes rule (2) then yields the conditioned quantum state $\rho(s_{m})|_{e_{m}}$ for gridpoint $s_{m}$. Intuitively, function $f_{m}$ can be thought of as an affinity function on state space $\Omega$ measuring how “similar” the given state $\omega_{m}$ is to the states included in our training dataset. Building operator $e_{m}$ and conditioning the prior state $\rho(s_{m})$ then yields the posterior state $\rho(s_{m})|_{e_{m}}$ that encodes the state space affinities captured by $f_{m}$. We often refer to the mapping $\omega_{m}\mapsto e_{m}$ as the feature map (as in Section 3.2) and to $f_{m}$ as the feature vector produced by kernel $\kappa_{c}$ from product state vector $\omega_{m}$.
With respect to our basis representation of $\rho(s_{m})$ by vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$ and effect $e_{m}$ by matrix $\bm{e}_{m}\in\mathbb{R}^{L\times L}$ the conditioning operation takes the form
$\underline{\rho}_{m}|_{e_{m}}=\frac{\sqrt{\bm{e}_{m}}\underline{\rho}_{m}}{\lVert\sqrt{\bm{e}_{m}}\underline{\rho}_{m}\rVert_{2}}.$
To avoid having to compute the square root of matrix $\bm{e}_{m}$ we modify the way we form effect $e_{m}$. We use function $f_{m}^{1/2}\in H_{\Omega,N}$ to form multiplication operator $\tilde{A}_{m}=M_{f_{m}}^{1/2}$ and define effect $\tilde{e}_{m}=(\operatorname{Proj}_{L}\tilde{A}_{m})^{2}=(\tilde{A}_{m}^{L})^{2}$. Conditioning with respect to $\tilde{e}_{m}$ now takes the form
$\underline{\rho}_{m}|_{\tilde{e}_{m}}=\frac{\sqrt{\bm{\tilde{e}}_{m}}\underline{\rho}_{m}}{\lVert\sqrt{\bm{\tilde{e}}_{m}}\underline{\rho}_{m}\rVert_{2}}=\frac{\bm{\tilde{A}}_{m}^{L}\underline{\rho}_{m}}{\lVert\bm{\tilde{A}}_{m}^{L}\underline{\rho}_{m}\rVert_{2}}.$
Although effects $\tilde{e}_{m}$ and $e_{m}$ are generally different, the two are equal when $L=N$ and thus also share the same asymptotic limit as $N\rightarrow\infty$ and $L\rightarrow\infty$ [55].
The data map $W_{J}$ embeds state vector $(x,s_{m})\in\Omega$ into $\mathbb{R}^{J}$ by considering the resolved state values at gridpoints neighboring $s_{m}\in S_{M}$, mimicking the structure of the stencil used for the spatial discretization of the considered dynamics. In this way kernel $k_{c}$ compares the stencil of points around $s_{m}$ to the corresponding stencil of points for every state included in our training data. This makes for a stronger affinity condition than comparing values of states at only a single gridpoint. For hyperbolic partial differential equations, where the notion of a local propagation velocity is well defined, using a stencil of values is analogous to using a time delayed signal.
Because the spatially local value $\omega_{m}=(x,s_{m})$ is compared to every other state vector present in our training data $\{\omega_{nm}\}_{nm=0}^{NM-1}$ spatial translations of state $x$ are naturally factored during the conditioning process. This means that if a spatial translation of state $x$ is present in our training dataset then the similarity between product state $\omega_{m}=(x,s_{m})$ and the corresponding value of the spatial translation of state $x$ will be identified in the conditioning process. For partial differential equations with spatial translation symmetry, this means that we do not have to artificially inflate our training dataset by including all spatial translations of the sampled states.
4.5 Multitrajectory data
To build the data driven formulation of our scheme we have so far relied on a single dynamical trajectory $x_{n}=\Phi^{n\Delta t}(x_{0})$, $x_{0}\in X$. In this section we generalize this construction to the case where our training data consists of multiple trajectories. We begin by motivating the use of multiple trajectories before presenting the modifications required in the theoretical and data driven formulations of the scheme.
Using a single dynamical trajectory is generally sufficient when considering a dynamical system possessing a unique physical measure, namely a measure that can be sampled by a set of initial conditions of positive measure with respect to the ambient measure [77]. These are measures whose sampling can be realized numerically, with SRB measures being an important example [78]. In such systems, a single long trajectory can be used to sample the physical measure of the dynamics.
Using multiple trajectories as training data is of interest when considering dynamical systems possessing multiple physical measures; for instance, hamiltonian systems or hyperbolic systems with several competing basins of attraction. Additionally, using multiple trajectories is necessary in the forecasting of parameter dependent systems, where one considers a dynamical system whose qualitative dynamics changes with respect to a bifurcation parameter.
Finally, even in systems with a unique physical measure, being able to use multiple trajectories to sample that measure allows one to consider optimal sampling strategies. In designing such a strategy one tries to optimize the sampling rate with respect to a chosen objective, instead of relying on the inherent rate provided by a single dynamical trajectory.
We consider a finite collection of invariant probability measures $\{\mu_{i}\}_{i}$ on $X$ with mutually disjoint supports where $i\in\{0,\ldots,I-1\}$ and $I\in\mathbb{N}$. We define the invariant probability measure $\mu=\sum_{i=0}^{I-1}\mu_{i}/I$ and form our quantum mechanical Hilbert space $H_{\Omega}=L^{2}(\Omega,\mu\times\nu_{M};\mathbb{C})$. Denoting $H_{\Omega,i}=L^{2}(\Omega,\mu_{i}\times\nu_{M};\mathbb{C})$ it follows that $H_{\Omega}$ admits the orthogonal decomposition
$H_{\Omega}\simeq\oplus_{i=0}^{I-1}H_{\Omega,i}$
[…]
To realize the above construction numerically we assume access to sampled trajectories $\{x_{n}^{(i)}\}_{n=0}^{N_{i}-1}$ of $N_{i}\in\mathbb{N}$ samples each, where $i\in\{0,\ldots,I-1\}$ as above. Similarly to the single trajectory case from Section 4.4, for each trajectory we define the empirical sampling measure $\mu_{i,N_{i}}=\sum_{n=0}^{N_{i}-1}\delta_{x_{n}^{(i)}}/N_{i}$ and assume that $\mu_{i,N_{i}}\to\mu_{i}$ in the weak-* sense as $N_{i}\to\infty$. We also define the empirical measure $\mu_{N}=\sum_{i=0}^{I-1}\mu_{i,N_{i}}/I$ and associated Hilbert space $H_{\Omega,N}=L^{2}(\Omega,\mu_{N}\times\nu_{M};\mathbb{C})$ with $N=\sum_{i=0}^{I-1}N_{i}$. By our assumption that the measures $\mu_{i}$ have mutually disjoint supports it follows that the sampled trajectories $\{x_{n}^{(i)}\}_{n=0}^{N_{i}-1}$ are also mutually disjoint. As a result we have the orthogonal decomposition
$H_{\Omega,N}\simeq\oplus_{i=0}^{I-1}H_{\Omega,N_{i}}$
with $H_{\Omega,N_{i}}=L^{2}(\Omega,\mu_{i,N_{i}}\times\nu_{M};\mathbb{C})$. Each component space $H_{\Omega,N_{i}}$ is a space of equivalence classes of functions represented by their values on the samples of the $i$th trajectory only. Leveraging the decomposition of $H_{\Omega,N}$ we use the approach described in Section 4.4 to build an eigenfunction basis for each component space $H_{\Omega,N_{i}}$ separately and use the $L_{i}\in\mathbb{N}$ leading eigenfunctions $\{\phi_{\ell}^{(i)}\}_{\ell=0}^{L_{i}-1}$ to form the subspace $H_{\Omega,N_{i}}^{L_{i}}\subset H_{\Omega,N_{i}}$. Denoting $L=\sum_{i=0}^{I-1}L_{i}$ we define the combined space
$H_{\Omega,N}^{L}=H_{\Omega,N_{0}}^{L_{0}}\oplus\cdots\oplus H_{\Omega,N_{I-1}}^{L_{I-1}}$
with the direct sum of the component bases producing a basis for the combined space $H_{\Omega,N}^{L}$.
The basis of $H_{\Omega,N}^{L}$ derived in this way consists of a total of $L$ elements, each being an eigenfunction of a block diagonal operator $K_{N}=K_{N_{0}}\oplus\ldots\oplus K_{N_{I-1}}$ with kernel integral operators $K_{N_{i}}\colon H_{\Omega,N_{i}}\to H_{\Omega,N_{i}}$ in its diagonal blocks. More specifically, each basis vector $\phi_{\ell}^{(i)}$ represents the values of a function on $N$ samples, with all of its values being zero except for the block of values that corresponds to the $N_{i}$ samples of the $i$th trajectory it was derived from. This approach has lower computational cost compared to building an eigenfunction basis for $H_{\Omega,N}$ by using a single kernel integral operator, which would require solving an eigenvalue problem for an $N\times N$ matrix. However, our approach also requires that each invariant measure $\mu_{i}$ is individually adequately sampled, as opposed to adequate joint sampling of measure $\mu$. This can potentially limit the number of high quality basis functions that can be derived from a dataset of fixed size $N$, for instance if one chooses to use a large number of trajectories with a small number of samples $N_{i}$ for each one.
Having derived the basis for space $H_{\Omega,N}^{L}$ the construction of the rest of the scheme proceeds as in Section 4.4. The only additional difference is in the definition of the shift operator $P_{N}$ where time shifts are now applied only within each trajectory, and the first temporal sample of each trajectory is mapped to zero. Namely, we make sure not to shift samples across different trajectories.
After forming our QMCl closure model in $H_{\Omega,N}^{L}$ our goal is to use it to predict the dynamics of a chosen dynamical system for out of sample initial conditions; namely, initial conditions lying outside of the basins of the measures $\mu_{i}$ sampled in the training stage. Thanks to the fact that the feature maps used to form quantum effects (Section 4.4) are not restricted to our training dataset, our closure scheme can ingest information from the resolved variables and render estimates of the fluxes even for out of sample data. Whether these estimates are sufficiently accurate ultimately depends on whether the kernel used for the conditioning of quantum states can identify samples of our training dataset whose corresponding flux evolution is similar to that of out of sample initial conditions. In the next section we present numerical results offering empirical evidence that out of sample predictions can indeed be made using QMCl for the tested dynamical system and family of initial conditions.
5 Application
The shallow water equations govern the evolution of a fluid when its vertical length scale can be considered negligible compared to its horizontal length scale. The equations are derived from the conservation laws for mass and momentum. Various forms of the equations have applications in areas such as ocean fluid dynamics, often incorporating additional terms modeling the effects of bottom topography and Coriolis force [68, 69].
We consider the shallow water equations on the one dimensional spatial domain $S=[-L_{S}/2,L_{S}/2]$ with periodic boundary conditions. The system of equations governs the evolution of water height $h(t,\cdot)$ and momentum $q(t,\cdot)=h(t,\cdot)v(t,\cdot)$ with $v(t,\cdot)$ denoting the velocity field and $t\geq 0$ time. The scalar fields $h(t,\cdot)$ and $q(t,\cdot)$ are assumed to take values in a sufficiently regular subset of $L^{2}(S,\nu;\mathbb{R})$ for all $t\geq 0$. The governing equations read
$\partial_{t}h+\partial_{s}q$ =0=0= 0 (10)
$\partial_{t}q+\partial_{s}(q^{2}/h+1/2Fr^{-2}h^{2})$ =0=0= 0
subject to an appropriate initial condition. In the above $\partial_{t}$ and $\partial_{s}$ denote differentiation with respect to time and space respectively; $Fr$ denotes the flow’s Froude number, which is a dimensionless ratio of inertial over gravitational forces. Using the state variable $x=(h,q)\in X\subset L^{2}(S,\nu;\mathbb{R}^{2})$ and function $f(x)=(q,q^{2}/h+1/2Fr^{-2}h^{2})$ we rewrite (10) in the compact form
$\partial_{t}x+\partial_{s}f(x)=0.$ (11)
The dynamics governed by (11) is equivariant under the actions of the spatial translation group on $S$.
We define the total energy of the fluid as the sum of its kinetic and potential energies
[…]
for all $t\geq 0$. If one assumes that the height and velocity fields are differentiable for all $t\geq 0$, thereby excluding discontinuous solutions, then the total energy is conserved along solution trajectories, namely $dE(t)/dt=0$. In this case the state space can be decomposed into the dynamics invariant disjoint subsets
$\mathcal{E}_{c}=\{x\in X\colon E(t)=c\text{ for all $t\geq 0$}\}$
each one defined as the subset of constant energy $c\geq 0$.
In the formation of discontinuities, the internal energy of the fluid is increased across a discontinuity, accompanied by an increase in entropy. Since the above definition of total energy does not include internal energy, the total energy is decreased across a discontinuity [69]. In the present work we consider only differentiable solutions.
5.1 Discretization
We employ a finite volume spatial discretization of (11) using a uniform fine grid of $M_{f}\in\mathbb{N}$ cells with length $\Delta s=L/M_{f}$ each, centered at points $s_{m}\in S$, $m\in\{0,\ldots,M_{f}-1\}$. We denote $S_{f}=\{s_{m}\}_{m=0}^{M_{f}-1}\subset M$ and $C_{m}=[s_{m}-\Delta s/2,s_{m}+\Delta s/2]\subset S$, $m\in\{0,\ldots,M_{f}-1\}$ where each cell $C_{m}$ can also be represented by its center point $s_{m}\in S_{f}$. We define the cell averaged state variable
[…]
reusing the same state symbol $x$. For any $t\geq 0$ function $x(t,\cdot)$ can be represented by the vector $\underline{x}\in\mathbb{R}^{2M_{f}}$ with $x_{m}=x(t,s_{m})$, $s_{m}\in S_{f}$.
We derive the semidiscretized system of equations governing the evolution of the state variables
$\dot{x}_{m}=\frac{1}{\Delta s}(F_{m}-F_{m+1})$ (12)
using the local Lax-Friedrichs (LLF) numerical flux
$F_{m}=\frac{1}{2}[f(x_{m-1})+f(x_{m})]-\frac{\lambda_{m}}{2}(x_{m}-x_{m-1})$ (13)
with
$\lambda_{m}=\max\,\bigl{(}\lvert q_{m-1}/h_{m-1}\rvert+Fr^{-1}\sqrt{h_{m-1}},\,\lvert q_{m}/h_{m}\rvert+Fr^{-1}\sqrt{h_{m}}\bigr{)}.$ (14)
In the above, subscripts denote components of the associated vectors while $F_{m}$ denotes the numerical flux at the left face of cell $C_{m}$ [79]. Based on (13) and (14) we introduce the notation
$F_{m}=\mathcal{F}_{LLF}(x_{m-1},x_{m})$
which will be used later on to define the coarse grid fluxes.
We now form a coarse grid and associated coarse variables defined as local spatial averages of the fine variables $x_{m}$. In particular we form $M_{c}\in\mathbb{N}$ coarse grid cells by merging every $M_{s}\in\mathbb{N}$ fine cells; in this way, $M_{c}=M_{f}/M_{s}$ and each coarse cell has length $\Delta\hat{s}=M_{s}\Delta s$. We define the resulting coarse grid $S_{c}=\{\hat{s}_{m}\}_{m=0}^{M_{c}-1}\subset S$ with cells $\hat{C}_{m}=[\hat{s}_{m}-\Delta\hat{s}/2,\hat{s}_{m}+\Delta\hat{s}/2]\subset S$, $m\in\{0,\ldots,M_{c}-1\}$. The coarse grid state variable $\hat{x}$ is formed using spatial averages over each coarse cell
$\hat{x}(t,s_{m})=\frac{1}{M_{s}}\sum_{j=M_{s}m}^{M_{s}(m+1)-1}x_{j},\quad s_{m}\in S_{c}$
and is represented by the vector $\underline{\hat{x}}\in\mathbb{R}^{2M_{c}}$ with $\hat{x}_{m}=\hat{x}(t,s_{m})$, $s_{m}\in S_{c}$. For the evolution of the coarse state variable $\hat{x}$ we derive the system of equations
$\dot{\hat{x}}_{m}=\frac{1}{\Delta\hat{s}}[(\hat{F}_{m}+G_{m})-(\hat{F}_{m+1}+G_{m+1})]$ (15)
with coarse grid fluxes
$\hat{F}_{m}=\mathcal{F}_{LLF}(\hat{x}_{m-1},\hat{x}_{m})$ (16)
and subgrid fluxes
$G_{m}=F_{M_{s}m}-\hat{F}_{m}.$ (17)
The system (15) is equivariant under the actions of the discrete group of spatial translations on $S_{c}$.
By definition, the coarse grid fluxes (16) depend only on the coarse variables. On the contrary, the subgrid fluxes (17) depend on both the coarse and the original fine variables; in particular, they depend on the values of $x_{m}$ in the fine cells that border the interfaces of each coarse cell. The subgrid fluxes act as antidiffusive corrections to the coarse grid fluxes, increasing the accuracy of the discretization compared to one performed entirely on the coarse grid [80].
The temporal discretization of evolution equations (12) and (15) is carried out using the modified Euler method (two-stage Runge-Kutta), which has favorable structure preservation properties for the present problem [80]. In what follows we consider the coarse mesh variables as the resolved degrees of freedom and the discrete time version of (15) as the true resolved dynamics. For the system to be complete we need to provide a closure model that can infer the subgrid fluxes $G_{m}$ from the resolved variables $\hat{x}_{m}$, thereby removing the explicit dependence on the fine mesh variables $x_{m}$.
5.2 Closure model
To define the appropriate state space $X$ for the dynamics we begin by considering the measure space $(S,\mathcal{A},\nu)$ where $\mathcal{A}$ denotes the Borel $\sigma$-algebra on the spatial domain $S$. We form $\mathcal{A}_{f}\subset\mathcal{A}$ as the $\sigma$-algebra generated by the disjoint subsets $C_{m}\subset S$ that correspond to the fine discretization cells; analogously, we form $\mathcal{A}_{c}\subset\mathcal{A}$ as the $\sigma$-algebra generated by the coarse cells $\hat{C}_{m}\subset S$. We define measures $\nu_{f}$ and $\nu_{c}$ as the restrictions of $\nu$ to the $\sigma$-algebras $\mathcal{A}_{f}$ and $\mathcal{A}_{c}$ respectively.
The state space of the true spatiotemporal dynamics (11) is an infinite dimensional Hilbert space $X\subset L^{2}(S,\nu;\mathbb{R}^{2})$. To be able to simulate the “true” dynamics numerically we approximate the state space by $X\subset L^{2}(S,\nu_{f};\mathbb{R}^{2})$ where $X$ is now considered as a Hilbert subspace of the space of $L^{2}$ functions on the mesh produced by our fine discretization. Next we define the resolved state space $\mathcal{X}=L^{2}(S,\nu_{c};\mathbb{R}^{2})\cap X$ and decompose the full state space as $X=\mathcal{X}\oplus\mathcal{Y}$ with unresolved state space $\mathcal{Y}=\mathcal{X}^{\perp}$. In this way the resolved state space can be considered as a space of piecewise constant functions, each taking a constant value on each cell formed by the coarse discretization. The full dynamics $\Phi^{n\Delta t}\colon X\to X$ is now given by the discrete time version of (12) and the resolved dynamics $\Phi_{\mathcal{X}}^{n\Delta\hat{t}}\colon X\to\mathcal{X}$ by the discrete time version of (15) with $n\in\mathbb{N}$ and timesteps $\Delta t>0$ and $\Delta\hat{t}=M_{s}\Delta t$.
We define the product state space $\Omega=X\times S$ and associated Hilbert space of observables $H_{\Omega,N}=L^{2}(\Omega,\sigma_{N};\mathbb{C})$ with total number of samples $N\in\mathbb{N}$ and measures $\sigma_{N}=\mu_{N}\times\nu_{M}$ and $\nu_{M}=\nu_{c}$. For notational consistency with the previous sections, in the following we denote by $M=M_{c}$ the number of coarse grid cells and by $S_{M}=S_{c}\subset S$ the set of grid points representing the coarse grid cells. We introduce two flux functions $z_{j}\in H_{\Omega,N}$, $j\in\{0,1\}$ producing the fluxes for the height and momentum variables respectively, denoting their collection by $z=\{z_{j}\}_{j=0}^{1}$. For every state $x\in X$ the fluxes $z_{j}(x)$ are given by (17).
Using the full dynamics we generate $N$ samples of the resolved state $\{\hat{x}_{n}\}_{n=0}^{N-1}$ and fluxes $\{z(x_{n},\cdot)\}_{n=0}^{N-1}$ which will be used as our training data and may consist of one or more dynamical trajectories. Each sample $\hat{x}_{n}$ or $z(x_{n},\cdot)$ is represented by a vector in $\mathbb{R}^{2M}$ holding the values corresponding to the height $h$ and momentum $q$ variables in each coarse grid cell.
To form our closure model we define the surrogate unresolved state space $\tilde{\mathcal{Y}}=L^{2}(S,\nu_{M};B(H_{\Omega}))$ and two quantum observables $A_{j}\in B(H_{\Omega,N})$, $j\in\{0,1\}$. Observable $A_{0}$ will be used to produce the surrogate fluxes for the height $h$ variables and is formed as multiplication by the true flux function $z_{0}$; analogously, observable $A_{1}$ will produce the surrogate fluxes for the momentum $q$ variables and is defined as multiplication by $z_{1}$. With these definitions in place, the surrogate fluxes are computed using the two surrogate flux functions $\tilde{z}_{j}\colon\tilde{\mathcal{Y}}\to H_{S}$ with action $\tilde{z}_{0}(\rho)(s_{m})=\operatorname{tr}(\rho(s_{m})A_{0})$ and $\tilde{z}_{1}(\rho)(s_{m})=\operatorname{tr}(\rho(s_{m})A_{1})$ where $\rho\in\tilde{\mathcal{Y}}$ and $s_{m}\in S_{M}$.
In A we explain the sequence of steps used to build the data driven QMCl scheme and apply it to predict the dynamics based on a given initial condition.
5.3 Numerical results
We consider the shallow water equations (11) on the spatial domain $S=[-25,25]$ with Froude number $Fr=1/\sqrt{2g}$ and $g$ denoting the acceleration of gravity. We perform the spatial discretization using a fine grid of $M_{f}=1920$ cells and the temporal discretization using the midpoint Euler method and a timestep $\Delta t=0.1\Delta s$ where $\Delta s$ is the length of each fine cell.
Training
We define the one-parameter family of initial conditions
[…]
with $s\in S$, $L_{S}=50$, parameter $0\leq\delta\leq 1$, $h_{0}$ the initial condition for height and $q_{0}=h_{0}v_{0}$ the initial condition for momentum. The solutions generated by this family of initial conditions feature a number of wave fronts traveling in different directions and interacting with one another. Figures 2 and 3 present space-time heatmaps of the solutions that will be used to test the predictive performance of our closure model.
Figure 1: Space-time heatmaps depicting two eigenfunctions $\phi_{\ell}$ of the kernel integral operator $\bm{K}_{N}$ from (8) for each of the three trajectories used as our training data. The top row shows the third eigenfunction for each of the three trajectories; the bottom row the fourth eigenfunction for each trajectory. From left to right, the first column corresponds to the trajectory data obtained for $\delta=0$, the second to $\delta=0.5$ and the third to $\delta=1$.
Using the assumptions and methodology introduced in Section 4.5 for dealing with multitrajectory training data, we consider $\mu$ to be the measure generated by the full dynamics (12) restricted to the family of initial conditions (18) for $\delta\in[0,1]$. To sample this measure we use three trajectories generated using the values $\delta\in\{0,0.5,1\}$. For each trajectory we integrate the full dynamics (12) in time for 12,200 timesteps and then sample the next 3,260 timesteps. Each sample contains the values of functions $h$ and $q$ on the fine grid.
To derive the resolved state samples that will be used as our training data, we coarsen the model and sampled data. By averaging the sampled data over every $M_{s}=20$ fine cells we arrive at a coarse grid of $M=96$ cells of length $\Delta\hat{s}=20\Delta s$ each. Given that each cell is now 202020 times longer we also increase the sampling timestep by the same amount to set $\Delta\hat{t}=20\Delta t$. From the original 3,260 samples per trajectory generated on the fine grid we keep every 20th sample, leading to 163 samples per trajectory. For each one of those samples we spatially average the solution over every 202020 cells to derive our resolved state samples, namely the averaged values of $h$ and $q$ over the coarse grid. We then derive the subgrid flux training samples using (17).
[…]
The number of delays $Q$ was chosen by a combination of physical reasoning and trial and error to balance the effectiveness of the derived basis in representing the true flux samples and the cost associated with its computation. A useful rule of thumb is to select a delay window that is comparable to the timescale of the physical processes of interest. In our case, we see that the traveling wave fronts interact approximately every 30 timesteps (1.5$\approx$ 1.5≈ 1.5 time units) for the solution in Figure 2 and approximately every 50 timesteps (2.5$\approx$ 2.5≈ 2.5 time units) for the solution in Figure 3, with similar results for additional solutions not shown here. Using $Q=64$ is sufficient to ensure that at least one period of this interaction is captured in each delay embedded sample. Additionally, increasing the value of $Q$ further did not lead to noticeable differences in our results. The stencil size $J$ used for the conditioning of quantum states was chosen empirically. Increasing the value of $J$ further was shown to slightly improve the performance of the bayesian conditioning operation, but with an associated increase in computational cost.
To form the orthonormal basis $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ we must compute the eigenvectors of an $NM\times NM$ symmetric kernel matrix $\bm{K}_{N}$ representing the data driven integral operator (8). To reduce the computational cost of this operation we perform a partial Cholesky factorization of $\bm{K}_{N}$ to approximate it by a lower rank matrix $\tilde{\bm{K}}_{N}=\bm{F}\bm{F}^{T}$ with $\bm{F}\in\mathbb{R}^{NM\times r}$ and $r<NM$. We perform the factorization using rank parameter $r=L=6144$ and the randomized algorithm developed in [66]. We then use the approximate kernel matrix to perform its bistochastic normalization and solve the eigenvalue problem with reduced cost [67].
A selection of the resulting eigenfunctions is presented in Figure 1, where we show two eigenfunctions for each of the three trajectories used to build our training dataset. The shown eigenfunctions exhibit propagating fronts which resemble the main features of the true state and flux samples presented in Figures 2 and 3. More specifically, they are approximately constant on the characteristic lines of the propagating wave fronts, as expected from their symmetry invariance property. Similar features are present in the rest of the eigenfunctions not shown here. As explained in Sections 4.2 and 4.3 the strong dynamic relevance of each eigenfunction is a result of its spatial symmetry invariance and of the fact that generally it cannot be represented as the tensor product of a single pair of spatial and temporal modes. These properties allow us to compute a compressed representation of the true dynamics with a moderate number of eigenfunctions, enhancing the computational efficiency of the resulting closure scheme.
Prediction
To test the predictive performance of the QMCl scheme we use two initial conditions obtained from (18) for $\delta\in\{0.25,0.75\}$ which are trajectories not included in the training dataset. After integrating the true dynamics (12) for 12,200 timesteps we spatially average the found solutions to generate the initial conditions that will be used to initialize the resolved state in QMCl. Using those initial conditions we integrate in time the coarse dynamics (15) using timestep $\Delta\hat{t}$ and the subgrid fluxes predicted by QMCl. In what follows we present results obtained after integrating for a total of 120 timesteps while performing bayesian conditioning of the quantum states every 10 timesteps.
Figure 2: Space-time heatmaps for the trajectory with $\delta=0.25$. The top row shows the results for the resolved state variables; the bottom row for the subgrid fluxes. For each row, from left to right: the first two plots show the results for the height field $h$, with the true results on the left and the predicted results on the right; the next two plots show the results for the momentum field $q$, with true results on the left and predicted on the right.
Figure 3: Space-time heatmaps for the trajectory with $\delta=0.75$. The layout of the plots is as in Figure 2.
The results comparing the evolution under the true and surrogate dynamics are grouped into four figures. Figures 2 and 4 include the results for the trajectory generated by the initial condition obtained from (18) for $\delta=0.25$. Figures 3 and 5 showcase the results obtained for $\delta=0.75$.
Figure 4: Spatial profiles of the final solution snapshot for the trajectory with $\delta=0.25$. The top row shows the results for the resolved state and the subgrid fluxes for the height field $h$; the bottom row for the momentum field $q$. In each row, the true results are on the left subplot, the predicted ones on the right subplot. Within each subplot, full lines are used for the resolved state with their numerical values indicated on the left axis; dashed lines are used for the corresponding fluxes with their values indicated on the right axis.
Figure 5: Spatial profiles of the final solution snapshot for the trajectory with $\delta=0.75$. The layout of the plots is as in Figure 4.
Figure 6: Space-time heatmaps for the trajectory with $\delta=0.25$ comparing the dynamics on the coarse grid with the surrogate dynamics obtained by QMCl. From left to right, the first two plots show the results for the height field $h$, with the coarse grid results on the left and the QMCl results on the right; the next two plots show the results for the momentum field $q$, with coarse grid results on the left and QMCl on the right. The colorbars for each subplot are as in Figure 2.
Looking at the heatmaps for the height and momentum fields in Figure 2 for the trajectory with $\delta=0.25$ and in Figure 3 for the one with $\delta=0.75$, we see that QMCl clearly captures the main features of the dynamics for the tested time horizon. For both trajectories QMCl correctly predicts the main traveling wave features of the dynamics, including the spatial variations of the waves as they interact with one another. Moreover, the magnitude of the predicted fields is consistent with the true dynamics in most regions of the spatial domain. In addition to the results shown in these figures, we have also verified that QMCl can reproduce the dynamical trajectories included in its training dataset with good accuracy.
In the present closure model the subgrid fluxes play the role of antidiffusive corrections to the fluxes generated entirely on the coarse grid [80]. Their values are proportional to the spatial derivative of the associated field, taking large values when the height or momentum field is changing rapidly in space, and approximately zero value when the fields are constant in space. As a result, the contribution of the subgrid fluxes is critical in regions of large spatial variation of the height and momentum fields, counterbalancing the increased diffusion of the solution introduced by our coarsening of the original dynamics by spatial averaging. The strong spatial variation of the true subgrid fluxes renders the considered dynamics challenging to resolve and predict accurately by any closure scheme (Section 5.4). Typically, when the exact value of the subgrid fluxes is not predicted correctly by the employed closure model (derived by QMCl or another method), the surrogate evolution will be more diffusive than the true evolution.
The effect of diffusion can be seen in the heatmaps for the two trajectories. By comparing the true and predicted fields of height and momentum, we see that the surrogate dynamics leads to stronger diffusion of the solution in areas where there is strong spatial variation. The increased diffusion is seen in the plots as bleeding of color between different space-time regions, with the boundaries between those regions being less sharp than in the true dynamics. Similarly, the same effect can be seen by comparing the subgrid flux fields for height and momentum. In the true dynamics, the peaks of the fluxes are sharp and clearly separated from the valleys, with clearly visible waves traveling in opposite directions. In the predicted dynamics, although the waves are still predicted correctly, their peaks are not as sharp, with their values diffusing more strongly to the surrounding region.
The flux plots also clearly demonstrate the effect of bayesian conditioning, which for these results is performed every 101010 timesteps (0.5$\approx$ 0.5≈ 0.5 time units). One can see the discontinuities that occur every 101010 timesteps for the predicted fluxes over large regions of the spatial domain. These discontinuities represent corrections introduced to the predicted flux values every time conditioning of the quantum states is performed. We perform the bayesian conditioning every 101010 timesteps to reduce the computational cost of the simulation, since the conditioning is one of the most expensive operations in our closure scheme (Section 5.4). In general, the prediction accuracy increases with increased conditioning frequency. Based on our numerical experiments, conditioning the quantum states every 101010 timesteps strikes a good balance between prediction accuracy and computational cost for the considered dynamics.
Similar conclusions can be drawn from the spatial profiles shown in Figure 4 for the trajectory with $\delta=0.25$ and Figure 5 for that with $\delta=0.75$. In these plots one can see that the subgrid fluxes are proportional to the spatial derivative of the associated field, leading to sharp variations in space. The predicted dynamics preserves the main features of the evolved fields, but struggles to capture the sharp peaks of the subgrid fluxes, leading to stronger diffusion of the height and momentum fields in regions of strong spatial variation.
To further demonstrate the crucial role played by the subgrid fluxes in counterbalancing diffusion, in Figure 6 we compare the evolution of the QMCl surrogate dynamics with the true dynamics generated entirely on the coarse grid, namely with the subgrid fluxes (17) set to zero at every grid point. The comparison is carried out for the trajectory generated from the initial condition (18) with $\delta=0.25$ (Figures 2 and 4), but similar conclusions can be drawn from other initial conditions. The comparison demonstrates that the surrogate fluxes predicted by QMCl successfully counteract the diffusive effect of the spatial coarsening seen in the coarse grid simulation for the tested time horizon. Namely, although the subgrid fluxes predicted by QMCl are not as accurate as the true subgrid fluxes in regions of strong spatial variation of the solution, they are nevertheless effective in counterbalancing diffusion and thus improving the accuracy of the predicted dynamics.
5.4 Discussion
In the following paragraphs we discuss several aspects of the presented closure scheme and how they inform future research directions.
Computational aspects
Extending the QMCl framework to spatiotemporal dynamics comes with the challenge of increased computational cost for its implementation, primarily due to the increased size of the training datasets required to capture the main features of dynamics that evolves over both space and time. As such, one key area of focus for future work is reducing the cost of the numerical implementation of QMCl and improving its scaling with sampling parameter $N$, spatial grid size $M$ and spectral resolution parameter $L$.
In the offline stage the most expensive operation is computing the leading $L$ eigenfunctions of the $NM\times NM$ symmetric kernel matrix $\bm{K}_{N}$ representing the kernel integral operator (8). Even for moderate size parameters $N$ and $M$, solving such an eigenvalue problem can quickly become very expensive. As explained in Section 5.3, to reduce this computational cost when generating our numerical results we performed a partial Cholesky factorization of $\bm{K}_{N}$ to approximate it by a lower rank matrix using the randomized numerical algorithm developed in [66]. We then used the approximate kernel matrix to perform its bistochastic normalization and solve the resulting eigenvalue problem with reduced cost [67].
As an alternative we also tried using the random Fourier features (RFF) method to approximate $\bm{K}_{N}$ by a lower rank matrix [81]. However, we found that a relatively large number of Fourier features was required to obtain a satisfactory approximation of kernel matrix $\bm{K}_{N}$ compared to the rank parameter $r$ required for the partial Cholesky factorization. The poor approximations obtained using the RFF method led to issues in the downstream bistochastic normalization of the approximate kernel matrix, often encountering negative or zero row sums that prevented its normalization. In addition to our empirical results, the difference in the accuracy of approximations obtained using partial Cholesky and RFF is also supported by the slower convergence rate of the RFF method [66]. We intend to continue investigating effective ways of reducing the cost of computing the kernel eigenfunction basis.
In the online stage the most expensive operation is the quantum state conditioning; more specifically, building the quantum effect operator from the current resolved state. To form the quantum effect operator we use kernel (9) to compare the current resolved state vector to all state vectors present in our training dataset. This is the only online operation that scales with the size of the training dataset $NM$, and requires that we have the training data loaded in memory throughout the online stage.
One way of alleviating this issue is to use the RFF method to approximate the out of sample evaluation of kernel (9), thereby eliminating the dependence on the training data [81, 82]. In our case this approach requires keeping in memory one $L\times L$ matrix for each random Fourier feature used. Namely, using $r$ Fourier features requires holding in memory $rL^{2}$ floating point numbers of the employed precision. In our experiments we found that a large number of Fourier features was required to resolve the fine spatial features of the state vectors, rendering this memory cost requirement unsustainable.
Another approach is to form a reduced representation of the training dataset and use that to form the quantum effect operator. For instance one could choose a subset of the original training dataset and rely only on that subset to perform the quantum state conditioning. The formation of that subset can be carried out using the kernel sampling algorithm that underlies the partial Cholesky factorization algorithms [66]. Our goal is to continue investigating these and similar directions for improving the performance of the quantum state conditioning operation.
In QMCl the surrogate fluxes are computed as the expected value of an observable with respect to a given quantum state. As a result, to compute fluxes that vary strongly over the spatial domain one has to ensure that the used quantum states are very sharp: obtaining a clear maximum in some regions of product space $\Omega$ and being close to zero everywhere else, formally similar to a finite sum of Dirac delta measures. Working with such sharp quantum states poses several numerical challenges.
Producing accurate but very sharp quantum states requires a very precise choice of bandwidth parameter $\epsilon^{\prime}$ used in kernel (9) for the bayesian conditioning. Having such a selective kernel means that the resulting quantum state leads to highly accurate fluxes when the correct part of the training dataset is identified as similar to the current resolved state. However, it can lead to strongly inaccurate fluxes when regions of the training dataset are misidentified as similar, compared to using a less selective kernel that leads to wider averaging of the employed quantum observable.
Moreover, the renormalization of the quantum state after its bayesian conditioning can become a source of numerical instability when working with a very selective kernel. This can occur when no part of the training dataset is identified by the conditioning kernel as being similar to the current resolved state, leading to an approximately zero feature vector and conditioned state. More generally, obtaining an approximately zero conditioned quantum state can happen when the prior quantum state and the observation encoded in the feature vector have disjoint supports. In those cases, attempting to normalize the quantum state to unit norm can yield unpredictable results. Working with a wider, less selective kernel makes this scenario less likely to occur, albeit with the tradeoff of less informative conditioning and thus less accurate predicted fluxes.
Another way of addressing this issue is resetting the prior quantum state to an uninformative state and repeating the conditioning operation. This requires erasing the historical information accumulated in the prior state, thereby placing a higher level of trust on the dynamical relevance of the observation. Building a better theoretical understanding of the conditions under which such a scenario might occur and the computational ways in which it can be addressed is a topic of future work.
In addition to numerical instability, even if a very sharp quantum state can be consistently obtained by using a very selective kernel, there still remains the challenge of representing this quantum state with respect to the eigenfunction basis. Resolving such fine features usually requires a large number of eigenfunctions, leading to expensive simulations. Given that most QMCl operators are represented numerically as $L\times L$ matrices, the overall simulation cost scales with $L^{2}$.
The challenges involved in producing and working with a very selective kernel function motivate our use of a kernel bandwidth tuning algorithm [73] and of the introduction of variable bandwidth [76, 64]. These measures help us define a kernel function that is tailored to the characteristics of our training dataset, which in turn allows us to strike a reasonable compromise between numerical stability and sharpness of the kernel used for bayesian conditioning. As a result, we are able to construct a closure scheme that can predict the main features of the considered dynamics with reasonable computational cost and without numerical instabilities.
Finally, the successful application of QMCl relies on the precise calibration of kernel bandwidth $\epsilon$, number of delays $Q$ and conditioning stencil size $J$. Having efficient, reliable algorithms for selecting these parameters based on a training dataset is very important. The tuning of bandwidth parameter $\epsilon$ can be achieved using the algorithm developed in [73] or the median rule studied in [83]. When applicable, choosing the number of delays $Q$ can be informed by the dimension of the attracting state space region sampled by the dynamical trajectories generating the training data [84]. However, such dimension estimates are rarely available for real world datasets and their underlying dynamics. It would be interesting to consider computational algorithms that can estimate a suitable range of values for $Q$ based on the given training dataset, potentially also accounting for the presence of noise [85, 86, 87].
Theoretical aspects
One of the distinguishing features of the QMCl framework is that the classical dynamics is first embedded into the noncommutative setting of operator algebras used in quantum mechanics and only then discretized to finite dimension. Namely, instead of directly discretizing functions on Hilbert space $H$ to obtain functions on finite dimensional subspace $H^{L}$ we first map functions on $H$ to operators in $B(H)$ and then discretize them to obtain operators in finite dimensional $B(H^{L})$. This process renders the discretization employed by QMCl positivity preserving, with positive classical observables and probability density functions mapped to positive operators [55].
In addition to structure preserving properties, this discretization process embeds the discretized quantities in a space of larger dimension than direct discretization. More specifically, direct discretization maps real functions on $H$ to functions on the $L$-dimensional subspace $H^{L}$. The QMCl discretization maps real functions on $H$ to selfadjoint real operators in $B(H^{L})$ which is a subspace of dimension $\frac{1}{2}L(L+1)$. It is interesting to explore whether this difference in dimension along with the noncommutative nature of $B(H^{L})$ allow QMCl to encode more information about the original dynamics compared to closure schemes performed in $H^{L}$. Using tools from control theory [88] and quantum information theory [89] one should be able to quantify the differences between the two discretization approaches and their consequences for downstream closure schemes.
Another question concerns the out of sample performance of QMCl. In Sections 4.5 and 5.3 we considered training data generated by a finite collection of initial conditions without reference to an explicit process of generating these initial conditions. A natural extension of this scenario is to consider an invariant measure $\mu$ admitting a disintegration into conditional measures $\mu_{i}$ sampled from a probability distribution of initial conditions. Interesting questions in this direction are to characterize the behavior of our closure scheme as the number of sampled trajectories increases and derive associated out of sample error bounds.
6 Conclusion
We have developed a framework for the dynamical closure of spatiotemporal dynamics which combines elements of the quantum mechanical closure (QMCl) framework developed in [56, 57, 55] and the vector valued spectral analysis (VSA) technique of [65]. A key element of the present extended framework is that it encodes statistical information about unresolved degrees of freedom in the form of a field of quantum density operators on the spatial domain. This results in a flexible architecture where density operators are collocated with the discretization cells of the coarse model for the resolved dynamics. The density operators communicate with the resolved dynamics by providing the required flux terms through expectations of quantum observables and by receiving information through operator valued feature maps. Moreover, our approach inherits the symmetry factorization properties of VSA during the basis construction process and can also be used with training data consisting of multiple dynamical trajectories. As with the original QMCl work, the process of embedding classical dynamics into the noncommutative setting of quantum mechanics leads to positivity preserving discretizations of classical observables and probability density functions.
We have applied QMCl to a closure problem for the shallow water equations on a periodic one dimensional domain. The numerical results demonstrate that QMCl can effectively predict the dynamics of the out of sample initial conditions tested in this work, generating height and momentum fields that agree both qualitatively and quantitatively with the true data. As expected from a dynamical system that demonstrates strong spatial variations in the subgrid fluxes, the surrogate dynamics is more diffusive than the true dynamics generated by a high resolution reference model, and eventually struggles to fully capture the sharp spatial features of the true solution.
With that said, we believe that the results act as a successful proof of concept for the extension of QMCl to spatiotemporal dynamics, especially given the modest amount of data used for its training. Everything else kept the same, we expect that increasing the size of the training dataset $N$ or the spectral resolution $L$ will lead to improved prediction accuracy.
Future research directions stemming from this work include improving the computational cost of the data driven implementation of QMCl (particularly the cost of quantum state conditioning in the online stage), as well as the development of strategies for sampling initial conditions in the training stage and kernel tuning/learning. Additionally, we plan to explore applications of QMCl to systems exhibiting spatiotemporal chaos.
Acknowledgements
DG and JS acknowledge support from the U.S. Department of Energy under grant DE-SC0025101. CV was supported as a postdoctoral researcher from this grant. DG acknowledges support from the U.S. Department of Defense, Basic Research Office under Vannevar Bush Faculty Fellowship grant N00014-21-1-2946. DCF was supported as a PhD student from this grant. We thank Ilon Joseph, Michael Montgomery and Ilya Timofeyev for helpful conversations on the topics of this work.
Appendix A Data driven implementation
The process of building and using the developed QMCl scheme to simulate the surrogate dynamics consists of two stages: the offline and online stages. Below we outline the steps followed in each stage.
Offline stage
The offline stage consists of the operations needed to build the QMCl scheme using the training data $\{\hat{x}_{n}\}_{n=0}^{N-1}$ and $\{z(x_{n},\cdot)\}_{n=0}^{N-1}$. This involves the following sequence of steps.
  1. 1.
    Using the resolved state samples $\{\hat{x}_{n}\}_{n=0}^{N-1}$ we calibrate the bandwidth parameter $\epsilon$ for kernel (5) and perform its bistochastic normalization.
  2. 2.
    We use the samples $\{\hat{x}_{n}\}_{n=0}^{N-1}$ to compute the leading $L$ eigenfunctions $\{\phi_{\ell}\}_{\ell=0}^{L-1}$ of integral operator (8), defining subspace $H_{\Omega,N}^{L}$ as their linear span. The eigenfunction basis is represented by matrix $\bm{\Phi}\in\mathbb{R}^{NM\times L}$ holding one eigenfunction in each column.
  3. 3.
    We form the transfer operator $P_{N}^{L}$ and represent it by matrix $\bm{P}_{N}^{L}\in\mathbb{R}^{L\times L}$.
  4. 4.
    We use the flux samples $\{z(x_{n},\cdot)\}_{n=0}^{N-1}$ to form two quantum observables and project them to operators $A_{0}$, $A_{1}\in B(H_{\Omega,N}^{L})$ represented by the matrices $\bm{A}_{0}$, $\bm{A}_{1}\in\mathbb{R}^{L\times L}$. The first observable is built using the flux samples $\{z_{0}(x_{n},\cdot)\}_{n=0}^{N-1}$ corresponding to the height $h$ variables; the second using $\{z_{1}(x_{n},\cdot)\}_{n=0}^{N-1}$ corresponding to the $q$ variables.
  5. 5.
    We define the discrete field of quantum states $\rho\in L^{2}(S,\nu_{M};B(H_{\Omega,N}^{L}))$ with one state corresponding to each coarse grid cell. Each quantum state $\rho(s_{m})\in Q(H_{\Omega,N}^{L})\subset B(H_{\Omega,N}^{L})$, $s_{m}\in S_{M}$ is represented by a vector $\underline{\rho}_{m}\in\mathbb{R}^{L}$.
  6. 6.
    Using the resolved state samples $\{\hat{x}_{n}\}_{n=0}^{N-1}$ we calibrate the bandwidth parameter $\epsilon^{\prime}$ of kernel (9) and compute the bandwidth function needed for its variable bandwidth generalization.
Online stage
In the online stage we use the operators built in the offline stage to timestep the surrogate dynamics. We denote by $\tilde{x}\in\mathcal{X}$ the resolved state predicted by QMCl to distinguish it from the true resolved state $\hat{x}$. The online stage consists of the following steps.
  1. 1.
    We initialize the resolved state $\tilde{x}$.
  2. 2.
    We initialize the quantum states $\rho$ representing each grid cell using an uninformative initial condition: each quantum state $\rho(s_{m})$ is set to the projection along the unit function $1_{\Omega}\in H_{\Omega,N}^{L}$ that takes the value 111 everywhere.
  3. 3.
    We condition each quantum state using the initial resolved state $\tilde{x}$ and training data $\{\hat{x}_{n}\}_{n=0}^{N-1}$.
  4. 4.
    We use the quantum states $\rho$ and quantum observables to compute the surrogate fluxes $\tilde{z}(\rho)$ corresponding to two flux values for each grid cell.
  5. 5.
    We begin the timestepping loop.
    1. (a)
      We update the resolved state $\tilde{x}$ using the surrogate fluxes.
    2. (b)
      We update the quantum states $\rho$ using the transfer operator.
    3. (c)
      We condition the quantum states using the updated resolved state and training data samples.
    4. (d)
      We compute the updated surrogate fluxes $\tilde{z}(\rho)$.
In the timestepping loop as presented above, the quantum states are conditioned at every timestep. In our presented numerical results we have elected to condition the states less frequently, by only performing bayesian conditioning every $N_{r}$ timesteps with parameter $N_{r}\in\mathbb{N}$.